# The Born approximation for the fixed energy Calderón problem

### v1

@@ Abstract.
The Born approximation of a potential in the context of the Calderón inverse problem is an object that can be formally defined in terms of spectral data of the Dirichlet-to-Neumann map of the corresponding Schrödinger operator. In this article, we prove, in the case of radial potentials in the Euclidean ball and any fixed energy, that this Born approximation is well-defined as a compactly supported radial distribution, and that the Calderón problem can be reformulated as recovering a potential from its Born approximation. In addition, we show that the Born approximation depends locally on the potential and captures exactly its singularities, and that the functional that maps the Born approximation to the potential is Hölder continuous. We also prove that the Born approximation converges to the potential in the high-energy limit. Moreover, we give an explicit formula for the Fourier transform of the Born approximation at any fixed energy, and illustrate how it can be used as the basis of an accurate procedure to approximate a potential from its Dirichlet-to-Neumann map.
@@ 1.1. Outline of the article
The Calderón problem for a Schrödinger operator, also known as the Gel’fand-Calderón problem, asks for the reconstruction of a potential from the knowledge of certain boundary measurements of the solutions to the corresponding Schrödinger equation; see [Cal80, Cal06] and [Gel57]. In this article, we are interested in the fixed-energy version of this problem: given an open, bounded domain $\Omega\subset\mathbb{R}^{d}$ with $\mathcal{C}^{1}$ boundary, an energy $\kappa\in\mathbb{C}$ and a real-valued potential $q$, consider the elliptic boundary-value problem for the Helmholtz equation
@@
The Dirichlet-to-Neumann (DtN) map associates the boundary value $f$ with the normal derivative on $\partial\Omega$ of the corresponding solution $u$ to (1.1). Whenever $\kappa$ is not a Dirichlet eigenvalue of $-\Delta+q$, this produces a well-defined linear operator
@@
    Stability. Is the reconstruction process stable? That is, find a modulus of continuity for $\Phi^{\kappa}$. This is not possible in general, since $(\Phi^{\kappa})^{-1}$ is not continuous (see, for example, [Ale88, AC08]). This is, therefore, an ill-posed inverse problem. Nevertheless, one can ask for conditional stability when additional requirements of regularity and boundedness are imposed on the class of admissible potentials. [^1] Put in a more abstract setting, for a given compact set $K\subset\mathcal{X}$ one tries to compute the modulus of continuity of $(\Phi^{\kappa})^{-1}|_{\Phi^{\kappa}(K)}$, which exists since $\Phi^{\kappa}|_{K}$ is a uniformly continuous homeomorphism.
@@
The uniqueness for $d\geq 3$ was proved in [SU87], and later in [Nac88, Nov88] with particular emphasis on the case of fixed energy. The two-dimensional case was solved in  [Nac96, AP06] for conductivities and [Buk08, BIY15] for potentials. These results are proved using the notion of Complex Geometric Optics solutions (CGO) from [Fad65] ($d\geq 3$) or different families of exponentially growing solutions of the equations when $d=2$.
Conditional stability was proved in dimensions $d\geq 3$ by [Ale88] with a logarithmic-type modulus on continuity, which was shown to be optimal in [Man01] (see also [KRS21] for a more detailed account of this issue). Improved stability at fixed energy was established in [Nov11, Isa11, ILX20] in dimension three and [San13, San15] in two dimensions. The instability estimates of [Man01] have been improved in the $\kappa\neq 0$ case in [Isa13, KUW21].
@@
This approach was initiated in the context of the Calderón problem in [BCMM22, BCMM24], and is based on the notion of approximation in inverse scattering that can be traced back to the work of Born [Bor26]. It was successfully applied in [DMMN24] to the Gel’fand-Calderón problem at zero energy; and a systematic exposition of this approach, which in principle applies to any inverse problem, was given in [MM24]. In the present context, it can be described as follows. The map $\Phi^{\kappa}$ is Fréchet differentiable and satisfies $\Phi^{\kappa}(0)=0$; if for every $q\in\mathcal{X}$ it is possible to find $q^{\mathrm{B}}_{\kappa}\in\mathcal{B}$ in some space $\mathcal{B}$ of functions or distributions solving
@@
(1.5) ${\mathcal{X}}$${\Phi^{\kappa}(\mathcal{X})}$${\mathcal{B}}$$\scriptstyle{\Phi^{\kappa}_{\mathrm{B}}}$$\scriptstyle{\Phi^{\kappa}}$$\scriptstyle{\mathrm{d}\Phi^{\kappa}_{0}}$            ${q\;}$${\Lambda_{q,\kappa}-\Lambda_{0,\kappa}}$${q^{\mathrm{B}}_{\kappa}}$$\scriptstyle{\Phi_{\mathrm{B}}^{\kappa}}$$\scriptstyle{\Phi^{\kappa}}$$\scriptstyle{\mathrm{d}\Phi_{0}^{\kappa}}$
The $\mathrm{d}\Phi_{0}^{\kappa}$ map is continuous but its range is not closed, hence $(\mathrm{d}\Phi_{0}^{\kappa})^{-1}$ is discontinuous and only conditionally stable. The bad behavior of the inverse of $\Phi^{\kappa}$ is expected to be completely captured by the linear approximation $\mathrm{d}\Phi^{\kappa}_{0}$ and, therefore, the nonlinear map $\Phi_{\mathrm{B}}^{\kappa}=(\mathrm{d}\Phi_{0}^{\kappa})^{-1}\circ\Phi^{\kappa}$ and its inverse $(\Phi_{\mathrm{B}}^{\kappa})^{-1}$ are expected to be continuous.
The main difficulty in implementing this strategy lies in showing that there exists a solution of equation (1.4). The form of the solution, in case it exists, was obtained in [BCMM22] for the three-dimensional ball and zero energy: explicit representation formulas for $(\mathrm{d}\Phi_{0}^{0})^{-1}$ were obtained establishing a connection between the Gel’fand-Calderón problem and the classical Hausdorff moment problem, which was later extended to conductivities in [BCMM24]. In [DMMN24], $q^{\mathrm{B}}_{0}$ and the factorization (1.5) were shown to exist when $\mathcal{X}$ is a quite general class of radial potentials in the open unit ball $\mathbb{B}^{d}\subset\mathbb{R}^{d}$. In addition, in [DMMN24] $(\Phi^{0}_{\mathrm{B}})^{-1}$ was shown to be injective and Hölder continuous, locally from the boundary, and a rather detailed description of $q^{\mathrm{B}}_{0}$ was obtained (recovery of singularities and asymptotic behavior from the boundary), as well as an explicit algorithm for reconstructing $q$ from $q^{\mathrm{B}}_{0}$, and a partial characterization of the set of radial DtN maps. A key ingredient in obtaining these results were the tools introduced by Simon [Sim99] in the context of inverse spectral theory of Schrödinger operators on the half-line, especially the notion of $A$-amplitude. In [MM24], the authors show that in fact the $A$-amplitude coincides with the notion of Born approximation for this inverse one-dimensional spectral problem.
In this work, we continue these investigations and show that the factorization (1.5) exists for the Gel’fand-Calderón problem at fixed energy $\kappa\in\mathbb{R}$, still assuming radial symmetry on the potentials. Again, this relies on the analysis of a one-dimensional inverse spectral problem, the difficulty being that the corresponding Born approximation is no longer Simon’s $A$-amplitude. The existence of such a Born approximation, which we call the $A_{\kappa}$-amplitude, is an important part of this work. The main results of this article are:
@@
    Explicit description of $\mathrm{d}\Phi^{\kappa}_{0}$ and its inverse. We show that $\mathrm{d}\Phi^{\kappa}_{0}$ maps potentials to operators that are rotationally invariant, and whose eigenvalues are certain moments of the potential. Also, we give an explicit formula for $(\mathrm{d}\Phi^{\kappa}_{0})^{-1}$. This is presented in 1 and Proposition 2.4.
@@
    Existence of the Born approximation. This is stated in 2. Using the explicit formula for $(\mathrm{d}\Phi^{\kappa}_{0})^{-1}$ and tools of spectral theory of Schrödinger operators on the half-line, most particularly the existence of an $A_{\kappa}$-amplitude (Theorem 4.4), we prove that $q^{\mathrm{B}}_{\kappa}$, the Born approximation at energy $\kappa$, exists as a radial distribution. We also present explicit formulas for $q^{\mathrm{B}}_{\kappa}$.
@@
    Regularity of $q^{\mathrm{B}}_{\kappa}$. In 3 we prove that $q^{\mathrm{B}}_{\kappa}$ coincides with an integrable function outside the origin, that $q^{\mathrm{B}}_{\kappa}-q$ is one derivative smoother than $q$, and that $q^{\mathrm{B}}_{\kappa}=q$ at the boundary.
@@
  5. v)
    Injectivity of $\Phi_{\mathrm{B}}^{\kappa}$ and stability of $(\Phi_{\mathrm{B}}^{\kappa})^{-1}$. The map $\Phi_{\mathrm{B}}^{\kappa}:q\mapsto q^{\mathrm{B}}_{\kappa}$ is injective in a strong sense: two potentials coincide in a neighborhood of the boundary if and only if their Born approximations at energy $\kappa$ coincide in that same neighborhood. In addition, $(\Phi_{\mathrm{B}}^{\kappa})^{-1}$ is Hölder continuous with respect to the $L^{1}$ norm in the complement of any arbitrarily small ball centered in the origin. This is the scope of 5, which is derived from its one-dimensional analogue Theorem 4.5.
Furthermore, as already noted in [DMMN24] when $\kappa=0$, these results also give a partial characterization of the set of DtN maps for this problem. We note that the approach based on one-dimensional inverse spectral theory has been applied in the context of the Steklov problem for warped product manifolds in [DKN21, DHN23, DKN23]. In particular, the results in [DKN21] imply stability and uniqueness results for the radial Calderón problem, both for the conductivity and Schrödinger cases. We also mention that spectral theory methods had already been used in the context of the radial Calderón problem in the work [KV85, Section 6], and that they have been used to produce convergent reconstruction algorithms in [Syl92] in the 2-$d$ conductivity case.
@@ 1.2. Statement of the main results
Recall that the Dirichlet spectrum of $-\Delta$ on the ball $\mathbb{B}^{d}$ can be expressed in terms of the zeros of certain Bessel functions. The operator $\Lambda_{q,\kappa}$ is well defined as soon as $q\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p$ admissible, and $\kappa\in\mathbb{C}\setminus\operatorname{Spec}_{H^{1}_{0}}(-\Delta+q)$. Since $q$ is radial, the DtN operator commutes with the action of $\operatorname{SO}(d)$ on $\mathbb{S}^{d-1}$; as a consequence, its eigenfunctions are spherical harmonics (see Appendix B for the definition and properties of spherical harmonics). More precisely, if $\mathfrak{H}^{d}_{\ell}$ denotes the subspace of spherical harmonics of degree $\ell\in\mathbb{N}_{0}$ on $\mathbb{S}^{d-1}$, then
@@
The $\kappa$-moments appear in the expression of the Fréchet differential of $\Phi^{\kappa}$ at $q=0$. In Proposition 2.4 we prove that, as soon as $\kappa$ is not a Dirichlet eigenvalue of $-\Delta$ in the ball, $\mathrm{d}\Phi^{\kappa}_{0}(q)$ is a bounded operator on $L^{2}(\mathbb{S}^{d-1})$ for every $q\in L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$. These operators turn out to be invariant by the action of $\operatorname{SO}(d)$ and are characterized by their restriction on spherical harmonics, which is precisely given by
@@
Let $U\subseteq\mathbb{R}^{d}$ be an open set. We denote by $\mathcal{E}^{\prime}(U)$ the space of distributions with compact support in $U$, and $\mathcal{E}^{\prime}_{\mathrm{rad}}(U)\subset\mathcal{E}^{\prime}(U)$ the subset of radially symmetric distributions.
@@
where ,×$\left\langle\cdot,\cdot\right\rangle_{\mathcal{E}^{\prime}\times\mathcal{C}^{\infty}}$ denotes the duality pairing in $\mathcal{E}^{\prime}(\mathbb{R}^{d})$. This makes sense since $(\varphi_{\kappa,\ell})^{2}\in\mathcal{C}^{\infty}(\mathbb{R}^{d})$.
Our first result is an explicit formula to reconstruct a distribution from its moments. We define the Fourier transform of $f\in\mathcal{E}_{\mathrm{rad}}^{\prime}(\mathbb{R}^{d})$ as
@@
Let $\mathcal{P}_{\ell,d}:L^{2}(\mathbb{S}^{d-1})\longrightarrow\mathfrak{H}_{\ell}^{d}$ be the orthogonal projection onto $\mathfrak{H}^{d}_{\ell}$. We denote by $Z_{\ell,d}$, $\ell\in\mathbb{N}_{0}$, the zonal functions, which are characterized by
@@
The zonal functions are directly related to the Legendre polynomials, more details are given in Appendix B.
Theorem 1.
where the series are absolutely convergent.
This result is proved in Section 3. Note that taking the limit $\kappa\to 0$ in (1.15) recovers the representation for $\kappa=0$ that was obtained in [BCMM22, DMMN24], see (1.24). To state the next result, we define
The elements of $\mathcal{B}_{d}$ are radial distributions in $\mathbb{B}^{d}$ that coincide with $L^{1}$ functions in a neighborhood of the boundary of $\mathbb{B}^{d}$. This implies that the extension by zero of a distribution in $\mathcal{B}_{d}$ is an element of $\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{R}^{d})$. In this way, when convenient, $\mathcal{B}_{d}$ can be identified with a subset of $\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{R}^{d})$. We adopt this perspective in a few points in this work, for example, to define the moments and the Fourier transform of elements of $\mathcal{B}_{d}$ using (1.13) and (1.14).
Note that (1.11) allows us to extend $\mathrm{d}\Phi^{\kappa}_{0}$ to $\mathcal{B}_{d}$:
$\mathrm{d}\Phi^{\kappa}_{0}:\mathcal{B}_{d}\longrightarrow\mathcal{L}(L^{2}(\mathbb{S}^{d-1})),$
and 1 can be viewed as an explicit formula for $(\mathrm{d}\Phi^{\kappa}_{0})^{-1}|_{\mathrm{d}\Phi^{\kappa}_{0}(\mathcal{B}_{d})}$. In particular, the theorem implies that, as soon as $q^{\mathrm{B}}_{\kappa}\in\mathcal{B}_{d}$ exists, it is uniquely determined by (1.12) in $\mathcal{B}_{d}$. For shortness, we define
Our next result shows the existence of $q^{\mathrm{B}}_{\kappa}$ in $\mathcal{B}_{d}$.
@@
and $q^{\mathrm{B}}_{\kappa}-q$ belongs to $\mathcal{C}^{m+1}(\mathbb{B}^{d}\setminus\{0\})$ if $q\in\mathcal{C}^{m}(\mathbb{B}^{d}\setminus\{0\})$ for any $m\in\mathbb{N}_{0}$.
@@
The identity $q^{\mathrm{B}}_{\kappa}|_{\mathbb{B}^{d}\setminus\{0\}}=q^{s}_{\kappa}$ holds in the sense of distributions and shows that $q^{\mathrm{B}}_{\kappa}$ can be identified with a locally integrable function on $\mathbb{B}^{d}\setminus\{0\}$. In other words, $q^{\mathrm{B}}_{\kappa}$ is a regularization of the singular function $q^{s}_{\kappa}$ in the sense of [GS64]. Note that (1.20) implies that $q|_{\mathbb{S}^{d-1}}=q^{\mathrm{B}}_{\kappa}|_{\mathbb{S}^{d-1}}$ when $q$ has a well-defined trace on the boundary. See Proposition 4.6 for a more precise estimate of the pointwise behavior of $|q^{\mathrm{B}}_{\kappa}-q|$ on $\mathbb{B}^{d}\setminus\{0\}$.
The phenomenon of recovery of singularities proved in this result is illustrated in Fig. 1 and Fig. 3.
We now show that in the high-energy limit $\kappa\to-\infty$, the Born approximation at energy $\kappa$ recovers the potential.
@@
    The convergence in (1.22), which involves entire functions, takes place in Fourier space. At this moment, it is not clear that this convergence can be expressed in a natural way in physical space; however, this should be possible at least for some classes of potentials. Evidence for this is presented in Fig. 3 below.
@@
Our final result shows two important properties of the map $(\Phi_{\rm B}^{\kappa})^{-1}$: a strong local form of injectivity and Hölder continuity.
Theorem 5.
Let $10<b<$ 1, $\kappa\in\mathbb{R}$, and $p$ be admissible.
  1. i)
    $q_{1,\kappa}^{\mathrm{B}}(x)=q_{1,\kappa}^{\mathrm{B}}(x)\;a.e.\text{ in }\,\{b<|x|<1\}\iff q_{1}(x)=q_{2}(x)\;a.e.\text{ in }\,\{b<|x|<1\}.$
  2. ii)
    $\max_{j=1,2}\left\lVert q_{j}\right\rVert_{L^{p}(\{b<|x|<1\})}<K,\qquad\left\|q^{\mathrm{B}}_{1,\kappa}-q^{\mathrm{B}}_{2,\kappa}\right\|_{L^{1}(\{b<|x|<1\})}<\varepsilon(b,\kappa),$
    one has
    (1.23) $\left\|q_{1}-q_{2}\right\|_{L^{1}(\{b<|x|<1\})}<C(b,\kappa,K)\left(\left\|q^{\mathrm{B}}_{1,\kappa}-q^{\mathrm{B}}_{2,\kappa}\right\|_{L^{1}(\{b<|x|<1\})}\right)^{1/(p^{\prime}+1)},$
    where $p^{\prime}$ is the Hölder conjugate exponent of $p$.
The local uniqueness phenomenon proved in the first part of this result is illustrated in Fig. 4.
Theorems 2 to 5 are proved in Section 5. The analogues Theorems 2, 3 and 5 in the case $\kappa=0$ have been proved in [DMMN24]. These results can be recovered from the ones in this work in the limit $\kappa\to 0$. For instance, the classical Hausforff moments, which played the same role as the $\kappa$-moments here, can be obtained from $\sigma_{\ell}[q,\kappa]$ by

### v2

@@ Abstract.
The Born approximation of a potential in the context of the Calderón inverse problem is an object that can be formally defined in terms of spectral data of the Dirichlet-to-Neumann map of the corresponding Schrödinger operator. In this article, we prove, in the case of radial potentials in the Euclidean ball and any fixed energy, that the Born approximation is well-defined as a compactly supported radial distribution, and that the Calderón problem can be reformulated as recovering a potential from its Born approximation. In addition, we show that the Born approximation depends locally on the potential and captures exactly its singularities, and that the functional that maps the Born approximation to the potential is Hölder continuous. We also prove that the Born approximation converges to the potential in the high-energy limit. Moreover, we give an explicit formula for the Fourier transform of the Born approximation at any fixed energy, and illustrate how it can be used as the basis of an accurate procedure to approximate a potential from its Dirichlet-to-Neumann map.
@@ 1.1. Outline of the article
The Calderón problem for a Schrödinger operator, also known as the Gel’fand-Calderón problem, asks for the reconstruction of a potential from the knowledge of certain boundary measurements of the solutions to the corresponding Schrödinger equation; see [Cal80, Cal06] and [Gel57]. In this article, we are interested in the fixed-energy version of this problem: given an open, bounded domain $\Omega\subset\mathbb{R}^{d}$, $d\geq 2$, with $\mathcal{C}^{1}$ boundary, an energy $\kappa\in\mathbb{C}$ and a real-valued potential $q$, consider the elliptic boundary-value problem for the Helmholtz equation
@@
The Dirichlet-to-Neumann (DtN) map associates the boundary value $f$ with the normal derivative on $\partial\Omega$ of the corresponding solution $u$ to (1.1). Whenever $\kappa$ is not a Dirichlet eigenvalue of $\Delta+q$, this produces a well-defined linear operator
@@
    Stability. Is the reconstruction process stable? That is, find a modulus of continuity for $\Phi^{\kappa}$. This is not possible in general, since $(\Phi^{\kappa})^{-1}$ is not continuous (see, for example, [Ale88, AC08, FKR14]). This is, therefore, an ill-posed inverse problem. Nevertheless, one can ask for conditional stability when additional requirements of regularity and boundedness are imposed on the class of admissible potentials. [^1] Put in a more abstract setting, for a given compact set $K\subset\mathcal{X}$ one tries to compute the modulus of continuity of $(\Phi^{\kappa})^{-1}|_{\Phi^{\kappa}(K)}$, which exists since $\Phi^{\kappa}|_{K}$ is a uniformly continuous homeomorphism.
@@
The uniqueness for $d\geq 3$ was proved in [SU87], and later in [Nac88, Nov88] with particular emphasis on the case of fixed energy (see also [CR16, Hab15]). The two-dimensional case was solved in [Nac96, AP06] for conductivities and [Buk08, BIY15] for potentials. These results are proved using the notion of Complex Geometric Optics solutions (CGO) from [Fad65] ($d\geq 3$) or different families of exponentially growing solutions of the equations when $d=2$.
Conditional stability was proved in dimensions $d\geq 3$ by [Ale88] with a logarithmic-type modulus on continuity, which was shown to be optimal in [Man01] (see also [KRS21] for a more detailed account of this issue). In dimension 2 it was proved in [BBR01, CFR10] (see also [FP18]). Improved stability at fixed energy was established in [Nov11, Isa11, ILX20] in dimension three and [San13, San15] in two dimensions. The instability estimates of [Man01] have been improved in the $\kappa\neq 0$ case in [Isa13, KUW21].
@@
This approach was initiated in [BCMM22, BCMM24] in the context of the Calderón problem, and is based on the notion of approximation in inverse scattering that can be traced back to the work of Born [Bor26]. It was successfully applied in [DMMN24] to the Gel’fand-Calderón problem at zero energy; and a systematic exposition of this approach, which in principle applies to any inverse problem, was given in [MM24]. In the present context, it can be described as follows. The map $\Phi^{\kappa}$ is Fréchet differentiable and satisfies $\Phi^{\kappa}(0)=0$; if for every $q\in\mathcal{X}$ it is possible to find $q^{\mathrm{B}}_{\kappa}\in\mathcal{B}$ in some space $\mathcal{B}$ of functions or distributions solving
@@
[…]
The objective is to exploit this factorization to obtain uniqueness, stability, reconstruction and characterization results for the inverse problem. The main difficulty in implementing this strategy is to show that there exists a solution of equation (1.4). In this work we show that (1.4) can be solved and that the factorization (1.5) exists for the Gel’fand-Calderón problem at fixed energy $\kappa\in\mathbb{R}$, assuming radial symmetry on the potentials.
In the radial case, the existence of $q^{\mathrm{B}}_{\kappa}$ for $\kappa=0$ has already been proved in [DMMN24], a result that also implies a partial characterization of the set of DtN maps. A key step in the proof involves the tools introduced by Simon [Sim99] in the context of inverse spectral theory of Schrödinger operators on the half-line, especially the notion of $A$-amplitude. In [MM24], the authors show that, in fact, the $A$-amplitude coincides with the notion of Born approximation for this one-dimensional inverse spectral problem.
Here, the proof of the existence of $q^{\mathrm{B}}_{\kappa}$ relies again on the analysis of a one-dimensional inverse spectral problem, the difficulty being that the corresponding Born approximation is no longer Simon’s $A$-amplitude. The existence of such a Born approximation, which we call the $A_{\kappa}$-amplitude, is an important part of this work. The main results of this article are:
@@
    Explicit description of $\mathrm{d}\Phi^{\kappa}_{0}$ and its inverse. We show that $\mathrm{d}\Phi^{\kappa}_{0}$ maps potentials to operators that are rotationally invariant, and whose eigenvalues are certain moments of the potential. Also, we give an explicit formula for $(\mathrm{d}\Phi^{\kappa}_{0})^{-1}$. This is presented in Theorem 3.1 and Proposition 2.4.
@@
    Existence of the Born approximation. This is stated in 1. Using the explicit formula for $(\mathrm{d}\Phi^{\kappa}_{0})^{-1}$ and tools of spectral theory of Schrödinger operators on the half-line, most particularly the existence of an $A_{\kappa}$-amplitude (Theorem 4.4), we prove that $q^{\mathrm{B}}_{\kappa}$, the Born approximation at energy $\kappa$, exists as a radial distribution. We also present explicit formulas for $q^{\mathrm{B}}_{\kappa}$.
@@
    Regularity of $q^{\mathrm{B}}_{\kappa}$. In 2 we prove that $q^{\mathrm{B}}_{\kappa}$ coincides with an integrable function outside the origin, that $q^{\mathrm{B}}_{\kappa}-q$ is one derivative smoother than $q$, and that $q^{\mathrm{B}}_{\kappa}=q$ at the boundary.
@@
    Injectivity of $\Phi_{\mathrm{B}}^{\kappa}$ and stability of $(\Phi_{\mathrm{B}}^{\kappa})^{-1}$. The map $\Phi_{\mathrm{B}}^{\kappa}:q\mapsto q^{\mathrm{B}}_{\kappa}$ is injective in a strong sense: two potentials coincide in a neighborhood of the boundary if and only if their Born approximations at energy $\kappa$ coincide in that same neighborhood. In addition, $(\Phi_{\mathrm{B}}^{\kappa})^{-1}$ is Hölder continuous, under mild a priori conditions on the potentials, with respect to the $L^{1}$ norm in the complement of any arbitrarily small ball centered in the origin. This is the scope of 3, which is derived from its one-dimensional analogue Theorem 4.5.
  5. v)
@@
In (1.5), the maps $\mathrm{d}\Phi_{0}^{\kappa}$ and $\Phi^{\kappa}$ are continuous, but the ranges of each of them are not closed. Hence $(\mathrm{d}\Phi_{0}^{\kappa})^{-1}$ and $\Phi^{\kappa}$ are both discontinuous and only conditionally stable. Therefore, an important consequence of iv) is that the bad behavior of the inverse of $\Phi^{\kappa}$ is captured by the linear approximation $\mathrm{d}\Phi^{\kappa}_{0}$, since the non-linear map $(\Phi_{\mathrm{B}}^{\kappa})^{-1}$ is Hölder stable.
@@ 1.2. Statement of the main results
Recall that the Dirichlet spectrum of $\Delta$ on the ball $\mathbb{B}^{d}$ can be expressed in terms of the zeros of certain Bessel functions. The operator $\Lambda_{q,\kappa}$ is well defined as soon as $q\in L^{p}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$, with $p$ admissible, and $\kappa\in\mathbb{C}\setminus\operatorname{Spec}_{H^{1}_{0}}(-\Delta+q)$. Since $q$ is radial, the DtN operator commutes with the action of $\operatorname{SO}(d)$ on $\mathbb{S}^{d-1}$; as a consequence, its eigenfunctions are spherical harmonics (see Appendix B for the definition and properties of spherical harmonics). More precisely, if $\mathfrak{H}^{d}_{\ell}$ denotes the subspace of spherical harmonics of degree $\ell\in\mathbb{N}_{0}$ on $\mathbb{S}^{d-1}$, then
@@
The $\kappa$-moments appear in the expression of the Fréchet differential of $\Phi^{\kappa}$ at $q=0$. In Proposition 2.4 we prove that, as soon as $\kappa$ is not a Dirichlet eigenvalue of $\Delta$ in the ball, $\mathrm{d}\Phi^{\kappa}_{0}(q)$ is a bounded operator on $L^{2}(\mathbb{S}^{d-1})$ for every $q\in L^{\infty}_{\mathrm{rad}}(\mathbb{B}^{d},\mathbb{R})$. These operators turn out to be invariant by the action of $\operatorname{SO}(d)$ and are characterized by their restriction on spherical harmonics, which is precisely given by
@@
The existence of a solution for (1.12) is not guaranteed. In order to prove that a solution exists, we need to allow for the possibility of $q^{\mathrm{B}}_{\kappa}$ being a distribution. Let $U\subseteq\mathbb{R}^{d}$ be an open set. We denote by $\mathcal{E}^{\prime}(U)$ the space of distributions with compact support in $U$, and $\mathcal{E}^{\prime}_{\mathrm{rad}}(U)\subset\mathcal{E}^{\prime}(U)$ the subset of radially symmetric distributions.
@@
where $\left\langle\cdot,\cdot\right\rangle_{\mathcal{E}^{\prime}\times\mathcal{C}^{\infty}}$ denotes the duality pairing in $\mathcal{E}^{\prime}(\mathbb{R}^{d})$. This makes sense since $(\varphi_{\kappa,\ell})^{2}\in\mathcal{C}^{\infty}(\mathbb{R}^{d})$.
To state our first result, we define
The elements of $\mathcal{B}_{d}$ are radial distributions in $\mathbb{B}^{d}$ that coincide with $L^{1}$ functions in a neighborhood of the boundary of $\mathbb{B}^{d}$. This implies that the extension by zero of a distribution in $\mathcal{B}_{d}$ is an element of $\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{R}^{d})$. In this way, when convenient, $\mathcal{B}_{d}$ can be identified with a subset of $\mathcal{E}^{\prime}_{\mathrm{rad}}(\mathbb{R}^{d})$. We adopt this perspective in a few points in this work, for example, to define the moments and the Fourier transform of elements of $\mathcal{B}_{d}$ using (1.13) and (1.14). Note that (1.11) allows us to extend $\mathrm{d}\Phi^{\kappa}_{0}$ to $\mathcal{B}_{d}$.
Let
We show, for potentials $X_{p,\kappa}$ with $p$ satisfying (1.6), the existence of $q^{\mathrm{B}}_{\kappa}$ as an element of $\mathcal{B}_{d}$ and establish a formula for its Fourier transform that shows, in particular, that $q^{\mathrm{B}}_{\kappa}$ is uniquely determined by (1.12) in $\mathcal{B}_{d}$ (see also Theorem 3.1). The Fourier transform of $f\in\mathcal{E}_{\mathrm{rad}}^{\prime}(\mathbb{R}^{d})$ is defined as
@@
We denote by $Z_{\ell,d}$, $\ell\in\mathbb{N}_{0}$, the zonal functions, which are characterized in terms of $\mathcal{P}_{\ell,d}:L^{2}(\mathbb{S}^{d-1})\longrightarrow\mathfrak{H}_{\ell}^{d}$, the orthogonal projection onto $\mathfrak{H}^{d}_{\ell}$, by
@@
The zonal functions are directly related to the Legendre polynomials; more details are given in Appendix B.
@@
and $q^{\mathrm{B}}_{\kappa}-q$ belongs to $\mathcal{C}^{m+1}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ if $q\in\mathcal{C}^{m}(\overline{\mathbb{B}^{d}}\setminus\{0\})$ for any $m\in\mathbb{N}_{0}$.
@@
The identity $q^{\mathrm{B}}_{\kappa}|_{\mathbb{B}^{d}\setminus\{0\}}=q^{s}_{\kappa}$ holds in the sense of distributions and shows that $q^{\mathrm{B}}_{\kappa}$ can be identified with a locally integrable function on $\mathbb{B}^{d}\setminus\{0\}$. In other words, $q^{\mathrm{B}}_{\kappa}$ is a regularization of the singular function $q^{s}_{\kappa}$ in the sense of [GS64, Chapter 1]. Note that (1.19) implies that $q|_{\mathbb{S}^{d-1}}=q^{\mathrm{B}}_{\kappa}|_{\mathbb{S}^{d-1}}$ when $q$ has a well-defined trace on the boundary. See Proposition 4.6 for a more precise estimate of the pointwise behavior of $|q^{\mathrm{B}}_{\kappa}-q|$ on $\mathbb{B}^{d}\setminus\{0\}$.
We prove two important properties of the map $(\Phi_{\rm B}^{\kappa})^{-1}$: a strong local form of injectivity and Hölder continuity under uniform $L^{p}(\mathbb{B}^{d})$ bounds for the potentials.
Theorem 3.
Let $\kappa\in\mathbb{R}$, $p$ be admissible and $b\in(0,1)$. Let $U_{b}:=\{x\in\mathbb{B}^{d}:b<|x|<1\}$.
  1. i)
    $q_{1,\kappa}^{\mathrm{B}}(x)=q_{1,\kappa}^{\mathrm{B}}(x)\;a.e.\text{ in }\,U_{b}\;\iff\;q_{1}(x)=q_{2}(x)\;a.e.\text{ in }\,U_{b}.$
  2. ii)
    $\max_{j=1,2}\left\lVert q_{j}\right\rVert_{L^{p}(U_{b})}<K,\qquad\left\|q^{\mathrm{B}}_{1,\kappa}-q^{\mathrm{B}}_{2,\kappa}\right\|_{L^{1}(U_{b})}<\varepsilon(b,\kappa),$
    one has
    (1.20) $\left\|q_{1}-q_{2}\right\|_{L^{1}(U_{b})}<C(b,\kappa,K)\left(\left\|q^{\mathrm{B}}_{1,\kappa}-q^{\mathrm{B}}_{2,\kappa}\right\|_{L^{1}(U_{b})}\right)^{1/(p^{\prime}+1)},$
    where $p^{\prime}$ is the Hölder conjugate exponent of $p$.
We finally show that in the high-energy limit $\kappa\to$-$\infty$, the Born approximation at energy $\kappa$ recovers the potential.
@@
    The convergence in (1.22), which involves entire functions, takes place in Fourier space. At this moment, it is not clear that this convergence can be expressed in a natural way in physical space; however, this should be possible at least for some classes of potentials.
@@
The local uniqueness phenomenon proved in the first part of 3 is illustrated in Fig. 4, whereas the phenomenon of recovery of singularities in 2 is illustrated in Fig. 1, Fig. 3 and Fig. 4. Evidence for the convergence of the Born approximations to the potential in physical space is presented in Fig. 3 below.
Theorems 1 to 4 are proved in Section 5. The analogues of Theorems 1, 2 and 3 in the case $\kappa=0$ have been proved in [DMMN24]. These results can be recovered from the ones in this work in the limit $\kappa\to 0$. For instance, the classical Hausforff moments, which played the same role as the $\kappa$-moments here, can be obtained from $\sigma_{\ell}[q,\kappa]$ by
@@
We note that the approach based on one-dimensional inverse spectral theory has been applied in the context of the Steklov problem for warped product manifolds in [DKN21, DHN23, DKN24]. In particular, the results in [DKN21] imply stability and uniqueness results for the radial Calderón problem, both for the conductivity and Schrödinger cases. We also mention that spectral theory methods had already been used in the context of the radial Calderón problem in [KV85, Syl92].